OpenStudy (anonymous):

A student in an algebra course has test scored of 38 and 68. What score on the third test will give the student an average of 60

OpenStudy (anonymous):

38+68+x=180pts since 60 +60+60=180 x=74

OpenStudy (anonymous):

38 + 68 + X 60 ----------- = ------- 3 116 + X = 180 X = 74

OpenStudy (anonymous):

64 not 74 sorry : )

OpenStudy (anonymous):

thank you :)

OpenStudy (gw2011):

The answer is 74